OpenStudy (anonymous):

Simplify. 8 + 3 [3-(-1)6] a.)22 b.)14 c.)-1 d.)2

OpenStudy (anonymous):

iqot 22 is that correct ?

OpenStudy (anonymous):

From what you've written, the answer is actually 35. You need to enter the question in again.

OpenStudy (anonymous):

those are the ONLY answers listed on my test . . .

OpenStudy (anonymous):

Well, there's nothing else I can do. Sorry.

OpenStudy (anonymous):

which one shouLd ienter ?

OpenStudy (anonymous):

I have no clue. They're all wrong given what you've written. Pick anything.

OpenStudy (anonymous):

kk thanks for heLpinq me out ! : )

OpenStudy (anonymous):

np

OpenStudy (anonymous):

hey im sorry idid type it in wronq ! 8 + 3 [3-(-1)6] the 6 is a exponent

OpenStudy (anonymous):

iqot 2 .

OpenStudy (anonymous):

It's 14 then.

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

\[8+3(3-(-1)^6)=8+3(3-1)=8+3\times 2=14\]

OpenStudy (anonymous):

how did you qet 14 ?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

kk !